# XAI-CF – Examining the Role of Explainable Artificial Intelligence in Cyber Forensics

### v2

@@ 7.0.2 Bias Management
In an XAI-CF system, the pieces of evidence produced should be unbiased. A bias can occur in an XAI-CF system for different reasons: Biased training data, such as a manually labeled dataset by a forensics analyst, that can affect the credibility of the model’s output; Human infused bias when interacting with XAI-CF; Stale data, such as when out-of-date forensics datasets are used for training. Interested readers are referred to (mehrabi2021survey, ; schwartz2022towards, ) for a detailed list and discussion of AI biases. The problem of bias in AI is neither new nor unique and it is not possible to completely eliminate bias from an AI system. It is an ongoing process and there are many approaches for mitigating AI bias, such as training workshops that can employ procedures for unbiasing humans that interact with AI and the development of more diverse and comprehensive training data that addresses biases. Roselli et al. (roselli2019managing, ) recommend a set of processes to mitigate the impact of bias in AI systems. Schwartz et al. describe three major challenges of mitigating AI bias, datasets, testing, and human factors. The authors also provide recommendations to address these challenges. Stack et al. (slack2020fooling, ) provided a fooling method that created highly biased classifiers and was able to deceive state-of-the-art XAI techniques, such as LIME and SHAP in producing simple explanations that did not reflect the underlying biases in the system. There is a need to improve these state-of-the-art XAI techniques and in parallel develop new methods and approaches that can successfully manage the AI biases. Legal protection offered by non-discrimination laws is challenged when AI not humans discriminate (wachter2021fairness, ). There is also a need to devise laws, regulations, and standards that not only address human but also machine biases. We should also establish a continuous monitoring mechanism within the XAI-CF system to detect and address biases in real-time.
7.0.3 Adversarial Attacks
@@
XAI-CF can help an adversary without much effort to learn about AI forensic methods and plan new attack surfaces (anti-forensics) in a much better way. Is it possible and how to evade this situation? Adversarial attacks are one of the major challenges of an AI system (du2020sok, ; chakraborty2021survey, ). The adversary can manipulate the input to force the XAI-CF system to produce incorrect output. A pre-trained model used during an investigation loses its effectiveness and hence the pieces of evidence obtained lose their credibility. Different techniques are used by an adversary to attack an XAI system, such as data poisoning (baniecki2022fooling, ; aivodji2022fooling, ; baniecki2022manipulating, ), adversarial example (huang2023safari, ), and trust manipulation (lakkaraju2020fool, ), etc. A detailed list of attacks and their defenses are listed and discussed in (baniecki2023adversarial, ). Baniecki et al. (baniecki2023adversarial, ) reiterates in their survey about the apparent insecurities in XAI, that there are still several unaddressed attacks on explanation methods. To prepare and train an XAI-CF system against adversarial attacks one should know about the attack. Adversarial ML (huang2011adversarial, ) is a technique that learns from previous attacks by adversaries and plans new attacks that deteriorate the performance of the classifier. To counter these attacks different approaches are discussed in (martins2020adversarial, ) to adapt the classifiers to improve their classification. Generative Adversarial Networks (goodfellow2020generative, ) is another technique being applied to give insights into specific behaviors of adversaries to improve the defense against their attacks. The two techniques Adversarial ML and Generative Adversarial Networks can be used to improve the defenses against adversarial attacks. An adversarial attack can also be used as an anti-forensics technique (garfinkel2007anti, ). There are several research works carried out to detect anti-forensics activities (garfinkel2007anti, ; rekhis2011system, ; bhatt2017machine, ; hoelz2009artificial, ; mitchell2010use, ; sun2018novel, ; yu2016multi, ; chen2018densely, ; abozaid2023deployment, ; goel2023approach, ). To prevent anti-forensics activities, only one solution is provided and that is to preserve the privacy of the forensics evidence (armknecht2015privacy, ; nieto2016digital, ; nieto2017methodology, ) including digital files, emails, log files, and other documents. To develop counter-anti-forensics techniques is an open area of research (du2020sok, ).
7.0.4 Oversimplification
@@
The more accurate an AI system is the more difficult it is to explain and vice versa. One of the solutions to make AI explainable is to simplify the AI system. But there is a danger lurking in this process, we oversimplify the AI system to the extent that it may not produce the right results. Moreover, simplifying the explanation of a complex AI system, such as neural networks, will inevitably leave out important details that may hinder the discovery of new knowledge, i.e., the learning part of an AI system. To mitigate such risks we have to develop techniques that exchange accuracy with explainability in a controlled manner or provide explanation without affecting the accuracy. A lot of work has been done in the area of explainability in cognitive psychology (von2004explanation, ; keil2006explanation, ; schank2013explanation, ) but the same has not been systematically transferred to AI. Miller et al. (miller2019explanation, ) discuss in detail how we should develop methods and approaches to infuse this valuable research into XAI. Humans select explanations that are simpler than their full counterparts (miller2019explanation, ). Explanation of the XAI-CF system should be provided and simplified in a way that is to the satisfaction of the stakeholders of CF and also does not compromise the accuracy of the system. Thagard (thagard1989explanatory, ) gave seven foundational principles that explanation must comply with to be acceptable. Out of these seven, he contends the dependency of a better explanation on two principles simpler and more general. If all other things are equal, simpler explanations – those that cite fewer causes, and more general explanations – that explain more events are better explanations. Later this theory, that humans prefer simpler and more general explanations, is tested and proven true through an empirical study (read1993explanatory, ). In this context and as mentioned before, we recommend a simpler interactive visualization (e.g., a causal loop or stack and flow diagram) depicting more events that can provide a better explanation of an XAI-CF system.
7.0.5 CF and AI Chasm
@@
The disconnect between CF and AI communities (baggili2019founding, ) constitutes one of the major socio-technical challenges to developing a practical XAI-CF system. There is a lack of collaboration between these two communities that is not letting researchers from both these areas work together to bridge this gap. Moreover, the majority of the CF practitioners have limited knowledge of AI (sanchez2019practitioner, ; abraham2021automatically, ). For example, as mentioned before, one solution to improve HCI in an XAI-CF system is to transfer basic AI knowledge to CF personnel. Furthermore, including a human-in-the-loop (CF practitioner) complements an XAI-CF system. This is only possible if there is a strong collaboration between these two communities and CF practitioners and other stakeholders start learning the basics of AI. Currently, researchers have pointed out this lack of disconnect (baggili2019founding, ; sanchez2019practitioner, ; abraham2021automatically, ) between CF and AI but there is no concrete work done to bridge this gap. Firstly, we recommend conducting workshops that bring together researchers and practitioners from both CF and AI communities to discuss and raise awareness of this topic. We also need to include law enforcement agencies in such workshops. Secondly, we need to develop standards, notations, formalism of concepts, and explanations that are easy to understand by both communities. In this context, for easier transfer of results and information, there is a need to use systematic definitions and simpler and more general explanations of XAI systems. This way it will be easier for a CF practitioner to learn and interact with an XAI-CF system.
@@ 8 Threats to Validity
For conducting and completing an unbiased study and literature review, and to extract the information correctly we mostly used three reviewers for collecting the data. Because of the scarcity of resources, we used the same three reviewers for reviewing all the papers. There is no unified state-of-the-art approach or technique for comparing our work with others, therefore our comparison is only indicative and symbolic. To make the comparison and analysis of the research works presented in this paper as fair as possible and to mitigate the threats mentioned above, we developed and used the research methodology as discussed in section 2. Further, we used the principle of triangulation (yin2015qualitative, ; patton2014qualitative, ) to increase the credibility of the study carried out in this paper. The triangulation principle is the way of seeking at least three ways of verifying a procedure, piece of data, or finding. We used scite, an AI-powered research engine, Google Scholar search, and manual confirmation of the collected data from different sources, i.e., data triangulation. We used three reviewers to check and confirm the procedures and methods adapted for the study carried out in this paper, i.e., methodological triangulation. Besides all these efforts for objectivity, there may be chances of relativism (yin2015qualitative, ; def-relativism, ) present in the study carried out in this paper.
@@ 9 Conclusion
CF is facing many new challenges, such as the rise of complex cyber devices, the proliferation of operating systems and file formats, pervasive encryption, the use of the cloud for remote processing and storage, and legal standards. To overcome these challenges requires new techniques, such as from the field of AI. To apply these techniques successfully in CF, we need to justify and explain the results to the stakeholders of CF, such as forensic analysts and members of the court, for them to make an informed decision. If we want to apply AI successfully in CF, there is a need to develop trustworthy, authentic, interpretable, understandable, and interactive AI systems. An XAI-CF system can play this role but is indispensable and is still in its infancy. In this paper, we explore and make a case for the significance and advantages of XAI-CF. We strongly emphasize the need to build a successful and practical XAI-CF system and discuss some of the main requirements and prerequisites of such a system. We present a formal definition of the terms CF and XAI-CF and a comprehensive literature review of previous works that apply and utilize XAI to build and increase trust in CF. To make the reader familiar with the study carried out in this paper, in addition to background, we also present a critical and short review of the works carried out in the last ten years in XAI and CF. We discuss some challenges facing XAI-CF, such as adversarial attacks, bias management, oversimplification, CF and AI chasm, and human-computer interaction. We also provide some concrete solutions to these challenges. We identify key insights and future research directions for building XAI applications for CF. This paper is an effort to explore and familiarize the readers with the role of XAI applications in CF, and we believe that our work provides a promising basis for future researchers interested in XAI-CF.

### v3

@@ 7.2 Bias Management
In an XAI-CF system, the pieces of evidence produced should be unbiased. A bias can occur in an XAI-CF system for different reasons: Biased training data, such as a manually labeled dataset by a forensics analyst, that can affect the credibility of the model’s output; Human-infused bias when interacting with XAI-CF; Stale data, such as when out-of-date forensics datasets are used for training. Interested readers are referred to [220, 221] for a detailed list and discussion of AI biases. The problem of bias in AI is neither new nor unique, and it is not possible to eliminate bias from an AI system. It is an ongoing process, and there are many approaches for mitigating AI bias, such as training workshops that can employ procedures for unbiasing humans that interact with AI and the development of more diverse and comprehensive training data that addresses biases.
Hanna et al. [222] provide a foundational overview of the major ethical challenges associated with AI systems, with a strong emphasis on algorithmic bias and its implications across the AI lifecycle. Roselli et al. [223] recommend a set of processes to mitigate the impact of bias in AI systems. Schwartz et al. [221] describe three major challenges of mitigating AI bias: datasets, testing, and human factors. The authors also provide recommendations to address these challenges. Stack et al. [224] provided a fooling method that created highly biased classifiers and was able to deceive state-of-the-art XAI techniques, such as LIME and SHAP, in producing simple explanations that did not reflect the underlying biases in the system. There is a need to improve these state-of-the-art XAI techniques and, in parallel, develop new methods and approaches that can successfully manage the AI biases. Legal protection offered by non-discrimination laws is challenged when AI, not humans, discriminate [225]. There is also a need to devise laws, regulations, and standards that not only address human but also machine biases. We should also establish a continuous monitoring mechanism within the XAI-CF system to detect and address biases in real-time.
7.3 Adversarial Attacks
@@
XAI-CF can help an adversary without much effort to learn about AI forensic methods and plan new attack surfaces (anti-forensics) in a much better way. Is it possible, and how to evade this situation? Adversarial attacks are one of the major challenges of an AI system [4, 226]. The adversary can manipulate the input to force the XAI-CF system to produce incorrect output. A pre-trained model used during an investigation loses its effectiveness, and hence, the pieces of evidence obtained lose their credibility. Different techniques are used by an adversary to attack an XAI system, such as data poisoning [227, 228, 229], adversarial example [230], and trust manipulation [231], etc. A detailed list of attacks and their defenses is listed and discussed in [232]. The surveys provided in [232, 233] reiterate about the apparent insecurities in XAI, that there are still several unaddressed attacks on explanation methods.
To prepare and train an XAI-CF system against adversarial attacks, one should know about the attack. Adversarial ML [234] is a technique that learns from previous attacks by adversaries and plans new attacks that deteriorate the performance of the classifier. To counter these attacks, different approaches are discussed in [235] to adapt the classifiers to improve their classification. Generative Adversarial Networks [236] is another technique being applied to give insights into specific behaviors of adversaries to improve the defense against their attacks. The two techniques, Adversarial ML and Generative Adversarial Networks, can be used to improve the defenses against adversarial attacks. An adversarial attack can also be used as an anti-forensics technique [237]. There are several research works carried out to detect anti-forensics activities [237, 238, 239, 240, 241, 242, 243, 244, 245, 246]. To prevent anti-forensics activities, only one solution is provided, and that is to preserve the privacy of the forensics evidence [247, 248, 249], including digital files, emails, log files, and other documents. To develop counter-anti-forensics techniques is an open area of research [4, 250].
7.4 Oversimplification
@@
The more accurate an AI system is, the more difficult it is to explain and vice versa. One of the solutions to make AI explainable is to simplify the AI system. But there is a danger lurking in this process; we oversimplify the AI system to the extent that it may not produce the right results. Moreover, simplifying the explanation of a complex AI system, such as neural networks, will inevitably leave out important details that may hinder the discovery of new knowledge, i.e., the learning part of an AI system. To mitigate such risks, we have to develop techniques that exchange accuracy with explainability in a controlled manner or provide explanations without affecting the accuracy. A lot of work has been done in the area of explainability in cognitive psychology [251, 252, 253], but the same has not been systematically transferred to AI. Miller et al. [254] discuss in detail how we should develop methods and approaches to infuse this valuable research into XAI. Humans select explanations that are simpler than their full counterparts [254]. Explanation of the XAI-CF system should be provided and simplified in a way that is to the satisfaction of the stakeholders of CF and also does not compromise the accuracy of the system. Thagard [255] gave seven foundational principles that explanation must comply with to be acceptable. Out of these seven, he contends that the dependency of a better explanation on two principles is simpler and more general. If all other things are equal, simpler explanations – those that cite fewer causes, and more general explanations – that explain more events are better. Later, this theory, that humans prefer simpler and more general explanations, is tested and proven true through an empirical study [256].
The key lies in shifting explainability from static, post-hoc visualizations toward dynamic, user-driven interaction, where users can probe, query, and manipulate model explanations in real time [257]. In this context and as mentioned before, we recommend a simpler interactive visualization (e.g., a causal loop or stack and flow diagram) depicting more events that can provide a better explanation of an XAI-CF system.
7.5 CF and AI Chasm
@@
The disconnect between CF and AI communities [258] constitutes one of the major socio-technical challenges to developing a practical XAI-CF system. There is a lack of collaboration between these two communities that is not letting researchers from both these areas work together to bridge this gap. Moreover, the majority of the CF practitioners have limited knowledge of AI [259, 260, 261]. For example, as mentioned before, one solution to improve HCI in an XAI-CF system is to transfer basic AI knowledge to CF personnel. Furthermore, including a human-in-the-loop [262] (CF practitioner) complements an XAI-CF system. This is only possible if there is a strong collaboration between these two communities, and CF practitioners and other stakeholders start learning the basics of AI.
Currently, researchers have pointed out this lack of disconnect [258, 259, 260] between CF and AI, but there is no concrete work done to bridge this gap. Firstly, we recommend conducting workshops that bring together researchers and practitioners from both CF and AI communities to discuss and raise awareness of this topic. We also need to include law enforcement agencies in such workshops. Secondly, we need to develop standards, notations, formalisms of concepts, and explanations that are easy to understand by both communities. In this context, for easier transfer of results and information, there is a need to use systematic definitions and simpler and more general explanations of XAI systems. This way it will be easier for a CF practitioner to learn and interact with an XAI-CF system.
7.6 Towards an XAI-CF System: A Conceptual Framework
Existing studies rarely integrate explainability, legal defensibility, and human-AI collaboration in a unified architecture. To bridge this gap, we propose a five-layer framework that embeds XAI across the entire forensic lifecycle. Figure 11 provides a high-level overview of the proposed conceptual framework. The framework employs a layered architecture consisting of five interconnected layers, with each layer relying on the output of the one before it.
Figure 11: Towards an XAI-CF system: A layered architecture.
The proposed XAI-CF framework uniquely integrates XAI principles throughout the entire digital forensic lifecycle. Unlike existing works that apply AI primarily for detection or classification, this framework:
  • 1.
    Enhances transparency and interpretability of all AI-driven steps.
  • 2.
    Supports human–AI collaboration during investigation.
  • 3.
    Aligns forensic outputs with legal admissibility requirements.
As such, it provides the first holistic blueprint for embedding explainability into modern CF practice, while leaving the detailed ontology for future work.
@@ 8 Threats to Validity
For conducting and completing an unbiased study and literature review, and to extract the information correctly, we mostly used three reviewers for collecting the data. Because of the scarcity of resources, we used the same three reviewers to review all the papers. There is no unified state-of-the-art approach or technique for comparing our work with others; therefore, our comparison is only indicative and symbolic. To make the comparison and analysis of the research works presented in this paper as fair as possible and to mitigate the threats mentioned above, we developed and used the research methodology as discussed in section 2. Further, we used the principle of triangulation [28, 29] to increase the credibility of the study carried out in this paper. The triangulation principle is the way of seeking at least three ways of verifying a procedure, piece of data, or finding. We used scite, an AI-powered research engine, Google Scholar search, and manual confirmation of the collected data from different sources, i.e., data triangulation. We used three reviewers to check and confirm the procedures and methods adapted for the study carried out in this paper, i.e., methodological triangulation. Besides all these efforts for objectivity, there may be chances of relativism [28, 263] present in the study carried out in this paper.
@@ 9 Conclusion
CF is facing many new challenges, such as the rise of complex cyber devices, the proliferation of operating systems and file formats, pervasive encryption, the use of the cloud for remote processing and storage, and legal standards. Overcoming these challenges requires new techniques, such as those from the field of AI. To apply these techniques successfully in CF, we need to justify and explain the results to the stakeholders of CF, such as forensic analysts and members of the court, for them to make an informed decision. If we want to apply AI successfully in CF, there is a need to develop trustworthy, authentic, interpretable, understandable, and interactive AI systems. An XAI-CF system can play this role, but it is indispensable and is still in its infancy. In this paper, we explore and make a case for the significance and advantages of XAI-CF. We strongly emphasize the need to build a successful and practical XAI-CF system and discuss some of the main requirements and prerequisites of such a system. We present a formal definition of the terms CF and XAI-CF and a comprehensive literature review of previous works that apply and utilize XAI to build and increase trust in CF. To make the reader familiar with the study carried out in this paper, in addition to the background, we also present a critical and short review of the works carried out in the last ten years in XAI and CF. We discuss some challenges facing XAI-CF, such as adversarial attacks, bias management, oversimplification, the CF and AI chasm, and human-computer interaction. We also provide some concrete solutions to these challenges. We identify key insights and future research directions for building XAI applications for CF. We introduce the first XAI-CF framework, designed to address the long-standing challenges of opacity, limited interpretability, and poor auditability in AI-assisted forensic workflows. This paper is an effort to explore and familiarize the readers with the role of XAI applications in CF, and we believe that our work provides a promising basis for future researchers interested in XAI-CF.